Graph each line using its y-intercept and slope
y + 4x = 8

Answer by emargo19(101)

Slope intercept equation (y=mx+b)
b = y-intercept
m=slope
y + 4x = 8
y = -4x + 8
b = 8 (Therefore y=8)
Lets plot y=8 and m= -4/1
I'll start by plotting y=8(.i., b=8)
From this point, I go down 4 on the y-axis from 8 and over 1 to the right on x-axis: